how much water must be evaporated from 65 gallons of a 9% salt solution to make a 12% salt solution?
Answer by addingup(3677)

0.09(65) = 0.12(65-x)
5.85 = 7.8-0.12x
-0.12x = -1.95
x = -1.95/-0.12 divide, and remember that -/- = + so you can ignore the - sign when you divide, since your answer will be positive anyway. I get that you have to evaporate 16.25 gallons.